OpenStudy (jhulian):

A ball is thrown straight down from the top of a 454-foot building with an initial velocity of -18 feet per second. Use the position function below for free-falling objects. s(t) = -16t2 + v0t + s0 What is its velocity after 3 seconds? v(3) = -114 ft/s What is its velocity after falling 100 feet? v =

OpenStudy (anonymous):

the vol after 3 second= 18 x 3

OpenStudy (whpalmer4):

no, velocity after 3 seconds is not 18 * 3

OpenStudy (jhulian):

i already got the V after 3s

OpenStudy (jhulian):

Its -114ft/s

OpenStudy (whpalmer4):

yes, agreed. now, how are you going to find the second part, any ideas?

OpenStudy (jhulian):

double differentiate?...lol...i have no idea

OpenStudy (whpalmer4):

if you know the time at which the position of the falling object is 100 feet below the starting point, can't you then calculate velocity?

OpenStudy (jhulian):

yeah .. you are right..but how do i get the time?

OpenStudy (whpalmer4):

what will the value of \(s(t)\) be at that point?

OpenStudy (jhulian):

8t^2+9t-177

OpenStudy (whpalmer4):

uh, no... what is \(s(0)=\)

OpenStudy (jhulian):

454

OpenStudy (whpalmer4):

we have \[s(t)=-16t^2+v_0(t)+454\] at the time of interest. the ball has fallen 10p feet from its position at \(s(0)\)

OpenStudy (whpalmer4):

that means \[s(t) = 454-100=354\]right?

OpenStudy (jhulian):

right

OpenStudy (whpalmer4):

so solve \[s(t)=354=-16t^2+v_0 t +454\]for \(t\)

OpenStudy (whpalmer4):

then plug the resulting (positive) value of \(t\) into the equation you used to calculate velocity.

OpenStudy (jhulian):

I got two ans: 5.27 and -4.14

OpenStudy (whpalmer4):

sounds about right. it is a quadratic, so two solutions, but only the positive one applies to this scenario.

OpenStudy (jhulian):

because time is not negative...right?

OpenStudy (jhulian):

so 5.27?

OpenStudy (whpalmer4):

yes, we are only using a portion of the parabola in our model of the position.

OpenStudy (whpalmer4):

and what is \{v(5.27)=\]

OpenStudy (jhulian):

i had -186.64 as my velocity

OpenStudy (whpalmer4):

sounds plausible. it accelerates at -32 ft/s^2 and has a bit more than 2 more seconds of that acceleration beyond your previous velocity of -114 ft/s.

OpenStudy (whpalmer4):

\[v_f=v_i+at\]

OpenStudy (whpalmer4):

any questions?

OpenStudy (jhulian):

lol... it says wrong ans...

OpenStudy (jhulian):

i put in 186.64 and then 187

OpenStudy (whpalmer4):

negative sign !

OpenStudy (jhulian):

sorry...i did put the negative sign

OpenStudy (jhulian):

like -186.64 and -187

OpenStudy (whpalmer4):

hang on a sec, checking your solution for \(t\)

OpenStudy (jhulian):

okay

OpenStudy (whpalmer4):

we have \[v_0=-18\text { ft/s}\]is that right?

OpenStudy (jhulian):

right

OpenStudy (whpalmer4):

if so, \[354=-16t^2-18t+454\]\[0=-16t^2-18t+100\]\[t=\frac{-(-18)\pm\sqrt{(-18)^2-4(-16)(100)}}{2(-16)}=\]\[\frac{18\pm\sqrt{6724}}{-32}=\frac{18\pm82}{-32}=2,-3.125\]

OpenStudy (jhulian):

hold up.. the time is 2s?

OpenStudy (whpalmer4):

yes, \[-16(2)^2-18(2)+45 =-64-36+454=354\]

OpenStudy (jhulian):

then the V is gonna be-82?...

OpenStudy (whpalmer4):

sorry, should have checked your value of \(t\) when y0u first got it, but it seemed like it was in the right ballpark.

OpenStudy (jhulian):

so then V= -32(t)-18 V= -32(2)-18 V=-82 Is that right?

OpenStudy (whpalmer4):

yes, just a sec, making a graph for you

OpenStudy (jhulian):

yeah...it is!!! Thanks so much! you are the best!

OpenStudy (whpalmer4):

Here you can see velocity and position changing together, with crosshairs at our two points of interest.

OpenStudy (jhulian):

Oh yeah...makes much sense now...

OpenStudy (whpalmer4):

a better one, forgot one of the lines:

OpenStudy (jhulian):

Once again thank you very much! you are a lifesaver

OpenStudy (whpalmer4):

Glad I could help. These problems are fun once you get the hang of them! Of course, neglecting air resistance, the answers are maybe not quite realistic all the time, but hey...

OpenStudy (jhulian):

yeah..lol I am gonna go off now. Do have a nice day!

OpenStudy (whpalmer4):

And you as well. Time for me to go fix the sweetie's burned out headlight...